john is x years old. his father is 3 times as old as he is. his mother is 3 years younger than his father. find the sum of their ages in terms of x in 10 years time.
Answer by ptaylor(2198)

NOW:
x=John's age
3x=father's age
3x-3=mother's age
TEN YEARS FROM NOW:
x+10=John's age
3x+10=father's age
(3x-3)+10=mother's age
SO THE SUM OF THEIR AGES IN TEN YEARS TIME IS:
x+10+3x+10+(3x-3)+10 get rid of parens
x+10+3x+10+3x-3+10 collect like terms
7x+27---------------ans

Hope this helps---ptaylor